OpenStudy (anonymous):

someone help me out with this easy proportional formula. i cant remember it!! I have the following number: 17,199.53. I need to transfer this amount into 2 pieces 1. the principle 2. 17.8% of the principle. I figured out that the correct numbers are 14,600.62 & 2,598.92 (which is 17.8% of 14,600.02). How can I structure this proportion into a formula so that I can multiply my principle by 17.8% and still equal the total

OpenStudy (anonymous):

Sorry, I thought I answered this 17,199.53=P(100+17.8)/100 P = 17,199.53 * 100 / (100+17.8)

OpenStudy (anonymous):

thank you - or i guess it can be simply stated as X/1.178 * (.178)

OpenStudy (anonymous):

thank you - or i guess it can be simply stated as X/1.178 * (.178)

OpenStudy (anonymous):

Hm

OpenStudy (anonymous):

Oh, yes. I see what you mean. that's the other (smaller) part.